OpenStudy (anonymous):

Evaluate the integral using integration by parts where possible. t^(1/3)lnt dt

OpenStudy (turingtest):

\[u=\ln t\]\[dv=t^{1/3}\]\[\int udv=uv-\intvdu\]try it

OpenStudy (turingtest):

*\[\int udv=uv-\int vdu\]try it

OpenStudy (anonymous):

dt = 1/t du right?

OpenStudy (turingtest):

du=(1/t)dt

OpenStudy (anonymous):

and v = (t^4/3)/(4/3)?

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

Okay, I got to (ln t)((t^4/3)/(4/3)) - I((t^4/3)/(4/3))*t^1/3 Where I = Integral. Now I'm lost. The Integral of ((t^4/3)/(4/3)) would be fraction on top of fraction..is that supposed to happen?

OpenStudy (turingtest):

gotta remember a little algebra sometimes ;)\[\frac1{4/3}=3/4\]so\[\int t^{1/3}\ln t=\frac34t\ln t^{4/3}-\frac34\int t^{4/3}\frac1tdt\]